Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 561
Entry time: 30 Oct 2024, 10:14
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 30.10.2024 |
|
|
General Announcement:
- Friday Nov. 1st is the deadline to submit the winter maintenance task list to XO.
CTRL:
- Nothing to report.
DA:
- Nothing to report.
EEE:
- Nothing to report.
VAC:
- Yesterday issue with SA3/PES in XTD10: during ZZ a leak developed and was vented by accident. Recovered during the day.
XPD:
- PES problem; thanks to VAC for support yesterday.
- XGM is again operational in SA3.
- A. Galler informs an 8D report will be done on the PES issue.
DET:
- Follow-up on MID AGIPD: it was tested and the issue did not appear again after power restart. It will be monitored in the coming days.
- Issue with AGIPD at HED (train IDs): fixed by workaround taking train IDs from DAQ. Seems to be related only to HED but with all AGIPD modules.
Needs to be followed up; S. Hauf adds it is likely to be the receiver board.
ITDM:
- Nothing to report.
XRO:
- Nothing to report.
LAS:
- Nothing to report.
UNSYS:
- Today a problem in SA3/cell25 happened with an Apple-X that does not move to Linear Polarization. Under investigation; Linear Vertical polarization will be anyway stablished with 3 segments.
- ZZ for the Phase shifter in SA3 is planned for Nov. 11th.
SPB:
- Nothing to report.
FXE:
- Last week stability issues reported: M. Knoll will follow-up.
MID:
- Nothing to report.
HED:
- Not present.
SCS:
- Not present.
SQS:
- Not present.
SXP:
- Not present. |